Solve the equation by first clearing the fractions or decimals.


51. 3.4n + 6.2 = -7.4
Any help with this equation would be Greatly appreciated.

Answer by checkley71(8403)

51) 3.4N+6.2=-7.4
34N+62=-74
34N=-74-62
34N=-136
N=-136/34
N=-4 ANSWER.
PROOF
3.4*-4+6.2=-7.4
-13.6+6.2=-7.4
-7.4=-7.4